on average there is one traffic accident in Chicago il during the 20 minutes of rush traffic. calculate the probability that there are at least three accident in the period from 9.00am to 11.00 am
Answer #1

Solution :

Given that ,

mean = = 20

Using poisson probability formula,

P(X = x) = (e- * x ) / x!

P(X 3 ) = 1 - P(X < 3)

= 1 - ( P(X = 0 ) + P( X = 1 ) + P( X = 2 )

= 1 - ( (e-3 * 30) / 0! + (e-3 * 31) / 1! + (e-3 * 32) / 2! )

= 1 - ( 0)

= 1

Probability = 1

The probability that there are at least three accident in the period from 9.00 am to 11.00 am is 1.